1)

A business incurs the following costs: Labor: $160/unit Materials: $40/unit Rent: $400,000/month Assume the firm produces 3 million units per month. The total variable cost, per month, is$ million. The total fixed cost, per month, is $ million The total cost is $ million.

2)

A local coffee shop is hoping to make use of its excess restaurant capacity in the evenings by experimenting with selling beer and wine. It speculates that the only additional costs are hiring more of the same sort of workers to cover the additional hours and costs of the new line of beverages. Which of the following are examples of hidden costs that are likely to emerge from this decision? Check all that apply Training costs for new and existing employees on beer and wine serving procedures Increased insurance premiums due to the presence of alcohol at the coffee shop The cost of maintaining the machines that make the coffee The monthly rent paid for the building The forgone revenues that could be earned by renting the coffee shop out for other events during evening hours 囗

3)

Students doing poorly in courses often consider dropping the courses. Many universities will offer a refund if the student drops a course before a deadline True or False: Prior to the deadline, students should not take this refund option into account when deciding to drop the course True False
Answer #1

1) The Total Cost is divided into two parts, one part depends on the level of output produced and is known as the Total Variable Cost, another part is independent of the level of output produced and is always fixed, this part is known as the Total Fixed Cost. Here, Rent is a part of Total Fixed Cost because irrespective of how much a producer produces, he has to pay a fixed amount of rent to his landlord at regular intervals of a certain period (say for example monthly, yearly etc). Now, Labor and Materials are a part of Total Variable Cost because the producer can choose to employ a fixed amount of labor and buy raw materials according to his needs and accordingly he can increase or decrease his purchases of materials and his demand for labor.

So, now according to the question assuming the firm produces 3 million units per month,

the Total Variable Cost, per month is = 3000000 * (Cost of Labor per unit + Cost of Materials per unit)

= 3000000 * (160 + 40)

= 3000000 * 200

= $ 600000000

= $ 600 million.

The Total Fixed Cost, per month is = $ 400000

= $ 0.4 million

The Total Cost is = Total Variable Cost + Total Fixed Cost

= 600 + 0.4   

= $ 600.4 million